# Fisher-Rao Gradient Flows of Linear Programs and State-Action Natural Policy Gradients

### v1

@@ 1. Introduction
Natural policy gradient (NPG) methods and their proximal and trust region formulations known as PPO and TRPO are among the most popular policy optimization techniques in modern reinforcement learning (RL). As such they serve as a cornerstone of many recent RL success stories including celebrated advancements in computer games [48, 49, 11] and the recent development of large language models like ChatGPT [1]. This has motivated a quickly growing body of work studying the theoretical aspects such as the convergence properties and statistical efficacy of natural policy gradient methods. Almost all of these works consider a specific model geometry where the Fisher-Rao metrics of the individual rows of the policy are mixed according to their state distribution or slight modifications of this [24, 9, 34, 26]. However, other choices for the model geometry are possible. In particular, the Fisher metric on the state-action distributions has been used to design a natural gradient method as well as actor-critic and a trust-region variant known as relative entropy search (REPS) [35, 36, 42]. This alternative natural policy gradient has been found to have the potential to reduce the severity of plateaus [35] and improve the performance of actor-critic methods [36].
Working with the state-action distributions has the significant benefit that the reward optimization problem becomes a linear program [25]. In particular, for rich enough parametric policy models, the state-action natural policy gradient flows follow the Fisher-Rao gradient flows of the state-action linear program [39]. This motivates us to study the convergence properties of Fisher-Rao gradient flows of general linear programs. These flows coincide with the solutions of entropy-regularized linear programs and thus by studying the convergence of the flow we also bound the error introduced by entropic regularization in linear programming.
@@ 1.1. Contributions
We provide an explicit convergence analysis of Fisher-Rao gradient flows of general linear programs and natural (policy) gradient flows with and without function approximation. More precisely, our main contributions can be summarized as follows:
  • We study Fisher-Rao flows of general linear programs and show linear convergence both in KL-divergence and in function value with an exponential rate depending on the geometry of the linear program, see Theorem 3.2.
  • In the case of non-unique optimizers, we show that the Fisher-Rao gradient flow converges to the information projection of the initial condition to the set of optimizers thereby characterizing its implicit bias, see Corollary 3.12.
@@
  • We study natural gradients for parametric measures, and show sublinear convergence under inexact gradient evaluations up to an approximation error and a distribution mismatch measured in the $\chi^{2}$-divergence, see Corollary 4.7.
@@ 1.2. Related works
State-action natural policy gradients were recently studied with and without state-action entropy regularization in [39]. For regularization strength $\lambda>0$ that work showed $O(e^{-\lambda t})$ convergence, but in the unregularized case the precise exponential rate was not characterized.
@@
In the case of a linear program, the Fisher-Rao gradient flow parametrized by time corresponds to the trajectory of solutions of the entropy-regularized program parametrized by the inverse regularization strength, which has been studied in several works. An exponential convergence result was obtained in [18] and subsequently the rate was characterized as $O(e^{-\delta t})$ for a constant $\delta$ depending on the geometry of the linear program [55, 51]. The results obtained in this article follow an alternative proof strategy and provide exponential convergence $O(e^{-\Delta t})$, where $\Delta\geq\delta$. This improvement can be strict for the linear programs encountered in Markov decision processes under standard assumptions. Whereas existing works study convergence in function value, our results also cover convergence in the KL-divergence. Finally, the geometry of Fisher-Rao gradient flows or equivalently the entropic central path was recently described as the intersection of the feasible region with a toric variety [50].
@@ 2. Preliminaries on Fisher-Rao Gradient Flows
Here, we provide a self-contained review of the properties of Fisher-Rao gradient flows that we require later. The results in this section can be generalized to a large class of Hessian geometries and – apart from the central path property – also to other objectives albeit with different proofs, for which we refer to [5, 37].
@@
where we denote the induced norm by $\lVert v\rVert_{g_{\mu}^{\operatorname*{\operatorname{FR}}}}\coloneqq{g_{\mu}^{\operatorname*{\operatorname{FR}}}(v,v)}^{\frac{1}{2}}$. The Fisher-Rao metric was introduced in the seminal works of C. R. Rao [45, 46] to provide lower bounds on the statistical error in parameter estimation known as the Cramer-Rao bound. This geometric approach to statistical estimation has subsequently led to the development of the field of information geometry, where N. N. Čencov characterized the Fisher-Rao metric as the unique Riemannian metric (up to scaling) that is invariant under sufficient statistics [17, 7, 8]. Despite its central role in statistics, our main motivation for studying the Fisher-Rao metric is for its use in reinforcement learning, where it has been used to design natural gradient algorithms as well as trust region methods [6, 35, 42]. Further, it is very closely related to entropic regularization in linear programming, which enjoys immense popularity particularly in computational optimal transport [44, 51], see also [55] for a detailed discussion of entropy regularized linear programming.
@@
as it is induced by the Hessian of the (negative) entropy, i.e., we have
$g_{\mu}^{\operatorname*{\operatorname{FR}}}(v,w)=v^{\top}\nabla^{2}\phi(\mu)w\quad\text{for all }v,w\in\mathbb{R}^{\mathbb{X}},\mu\in\mathbb{R}^{\mathbb{X}}_{>0}.$ (2.3)
As such, the Fisher-Rao metric falls into the class of Hessian metrics that have been studied in convex optimization; we refer to [5, 37] for general well-posedness and convergence results. An important concept in the analysis of Hessian gradient flows is the Bregman divergence induced by $\phi$, which in the case of the negative entropy is given by the KL-divergence
@@
Proposition 2.2 (Central path property).
@@
On the other hand, for the gradient flow we can use (2.8) and compute for $v\in TP$
$\langle\nabla\phi(\mu_{t})-\nabla\phi(\mu_{0}),v\rangle$ $=\int_{0}^{t}\partial_{s}\langle\nabla\phi(\mu_{s}),v\rangle{\mathop{}\!\mathrm{d}}s$
$=\int_{0}^{t}\langle\nabla^{2}\phi(\mu_{s})\partial_{s}\mu_{s},v\rangle{\mathop{}\!\mathrm{d}}s$
$=\int_{0}^{t}\langle\nabla f(\mu_{s})\mu_{s},v\rangle{\mathop{}\!\mathrm{d}}s$
$=\int_{0}^{t}\langle c,v\rangle{\mathop{}\!\mathrm{d}}s=t\langle c,v\rangle.$
@@
Corollary 2.3 (Sublinear convergence rate).
Consider Setting 2.1 and assume that the linear program (2.9) admits a solution $\mu^{\star}\in P$. Then it holds that
@@
By the central path property, we have
$c^{\top}\mu_{t}-t^{-1}D_{\operatorname{KL}}(\mu_{t},\mu)\geq c^{\top}\mu^{\star}-t^{-1}D_{\operatorname{KL}}(\mu^{\star},\mu).$
Rearranging yields the result. ∎
@@
Theorem 2.4 (Well-posedness of FR GFs).
@@
The local existence and uniqueness follows from the Picard-Lindelöf theorem. Hence, it suffices to show that the Fisher-Rao gradient flow does not hit the boundary $\partial P$ in finite time. By the central path property, this is equivalent to the statement that the solutions of all KL-regularized problems (2.10) lie in the interior $\operatorname{int}(P)$ of the polyhedron, which can be easily checked. ∎
@@ 3. Convergence of Fisher-Rao Gradient Flows
We have seen that Fisher-Rao gradient flows converge globally at a sublinear rate $O(t^{-1})$. We now build on this analysis and show that once the gradient flow enters a vicinity of the optimizer, it converges at a quasi-exponential rate $O(t^{\kappa}e^{-\Delta t})$, where $\Delta>0$ depends on the geometry of the linear program and $\kappa>0$ depends on the initial condition $\mu_{0}$. We consider linear programs of the following form.
@@
The following result is the main contribution of this article, where we defer the proof to Section 3.1. We first establish it under the assumption that the linear program (3.1) admits a unique solution, and provide a generalization in Theorem 3.13.
@@
where $\Delta>0$ and $t_{0}\geq 0$ are defined in (3.2) and (3.3), respectively, and
$R_{H}\coloneqq\max_{\mu\in P}H(\mu)-\min_{\mu\in P}H(\mu)\leq\log\lvert\mathbb{X}\rvert$ (3.9)
denotes the entropic radius of $P$.
@@
showing that Theorem 3.2 offers an improvement of these previous results. For the special case $P=\Delta_{\mathbb{X}}$, for which a matching lower bound was constructed in [55], the two constants agree. Indeed, it is easily checked that $\delta=\Delta$ if and only if there is a neighboring vertex $\mu\in N(\mu^{\star})$ which as minimal optimality gap $c^{\top}\mu^{\star}-c^{\top}\mu$ and has disjoint support from $\mu^{\star}$. For $P=\Delta_{\mathbb{X}}$ this is always true, but for Markov decision processes the feasible region of the (dual) linear program is a strict subset $\mathscrsfs{D}\subsetneq\Delta_{\mathbb{S}\times\mathbb{A}}$ and under the standard exploratory 5.2 we have $\delta<\Delta$. In Section 5.2 we provide an explicit example where $\delta<\Delta$.
Further, for gradient flows with respect to a Riemannian metric of the form
$g_{\mu}^{\sigma}(v,w)\coloneqq\sum_{x\in\mathbb{X}}\frac{v_{x}w_{x}}{\mu_{x}^{\sigma}}$ (3.11)
one can show $O(t^{-\frac{1}{\sigma-1}})$ convergence for $\sigma\in(1,2)$, see [39]. Note that this can be extended to the case $\sigma=2$, which corresponds to logarithmic barriers for which it is well known that the central path converges at a $O(t^{-1})$ rate [14, Section 11.2].
@@ 3.1. Convergence of Fisher-Rao gradient flows
Lemma 3.7.
Consider Setting 2.1, whereby we allow $f\colon\mathbb{R}_{>0}^{\mathbb{X}}\to\mathbb{R}$ to be an arbitrary differentiable function, and fix $\mu\in P$. Then for any $t\geq 0$ it holds that
@@
Using (3.15) and (3.12) we find that for $t\geq T$ it holds that
$\partial_{t}D_{\operatorname{KL}}(\mu^{\star},\mu_{t})=c^{\top}\mu_{t}-c^{\top}\mu^{\star}\leq-\kappa_{t}D_{\operatorname{KL}}(\mu^{\star},\mu_{t}).$
Now Gronwall’s inequality yields (3.13). By Corollary 2.3 we have for any $h>0$ that
$c^{\top}\mu^{\star}-c^{\top}\mu_{t}\leq\frac{D_{\operatorname{KL}}(\mu^{\star},\mu_{t-h})}{h}\leq D_{\operatorname{KL}}(\mu^{\star},\mu_{t_{0}})\cdot\frac{\exp\left(-\int_{t_{0}}^{t-h}\kappa_{s}{\mathop{}\!\mathrm{d}}s\right)}{h}.$
@@
Consider a polytope $P\subseteq\mathbb{R}^{\mathbb{X}}$ and let $c\in\mathbb{R}^{\mathbb{X}}$ be such that the linear function $\mu\mapsto c^{\top}\mu$ has a unique maximizer over $P$ at vertex $\mu^{\star}\in P$ and consider an arbitrary semi-norm $\lVert\cdot\rVert\colon\mathbb{R}^{\mathbb{X}}\to\mathbb{R}_{\geq 0}$. Setting $\frac{c}{0}\coloneqq+\infty$ for $c>0$, we have
$\Delta\coloneqq\min\left\{\frac{c^{\top}\mu^{\star}-c^{\top}\mu}{\lVert\mu^{\star}-\mu\rVert}:\mu\in N(\mu^{\star})\right\}>0,$ (3.16)
where $N(\mu^{\star})$ denotes the set of neighboring vertices of $\mu^{\star}$ and it holds that
$c^{\top}\mu^{\star}-c^{\top}\mu\geq\Delta\cdot\lVert\mu^{\star}-\mu\rVert\quad\text{for all }\mu\in P.$ (3.17)
@@
Note that since $\mu^{\star}$ is the unique maximizer of $\mu\mapsto c^{\top}\mu$ over $P$ it holds that $c^{\top}\mu^{\star}-c^{\top}\mu>0$ for every neighboring vertex $\mu\in N(\mu^{\star})$, which implies $\Delta>0$. If $\Delta=+\infty$, then $\lVert\mu^{\star}-\mu\rVert=0$ for all $\mu\in P$ and hence (3.17) holds trivially, so let $\Delta<+\infty$. The polytope $P$ is contained in the cone
$C=\left\{\mu^{\star}+\sum_{\nu\in N(\mu^{\star})}\alpha_{\nu}(\nu-\mu^{\star}):\alpha_{\nu}\geq 0\text{ for all }\nu\in N(\mu^{\star})\right\}$
generated by the edges adjacent to $\mu^{\star}$, see [59, Lemma 3.6]. Hence, for any $\mu\in P$ there are non-negative weights $\alpha_{\nu}\geq 0$ for $\nu\in N(\mu^{\star})$ such that
$\mu=\mu^{\star}+\sum_{\nu\in N(\mu^{\star})}\alpha_{\nu}(\nu-\mu^{\star}).$
Now we compute
$c^{\top}\mu^{\star}-c^{\top}\mu=\sum_{\nu\in N(\mu^{\star})}\alpha_{\nu}(c^{\top}\mu^{\star}-c^{\top}\nu)\geq\Delta\cdot\sum_{\nu\in N(\mu^{\star})}\alpha_{\nu}\lVert\mu^{\star}-y\rVert.$ (3.18)
Further, by the triangle inequality we have
$\lVert\mu^{\star}-\mu\rVert=\bigg{\lVert}\sum_{\nu\in N(\mu^{\star})}\alpha_{\nu}(\nu-\mu^{\star})\bigg{\rVert}\leq\sum_{\nu\in N(\mu^{\star})}\alpha_{\nu}\lVert\nu-\mu^{\star}\rVert,$
which completes the proof. ∎
@@
Consider a finite set $\mathbb{X}$ and a probability distribution $\mu\in\Delta_{\mathbb{X}}$. Let $c>1$ and set
$\delta\coloneqq\frac{c-1}{c+1}\cdot\min\{\mu_{x}:\mu_{x}>0\}>0.$ (3.19)
Then for all $\nu\in\Delta_{\mathbb{X}}$ satisfying $\lVert\mu-\nu\rVert_{\infty}\leq\delta$ it holds that
@@
$\begin{split}\mu_{x}\log\left(\frac{\mu_{x}}{\nu_{x}}\right)&=\mu_{x}\left(\log(\nu_{x}+(\mu_{x}-\nu_{x}))-\log(\nu_{x})\right)\\ &\leq\mu_{x}\left(\log(\nu_{x})+\frac{\mu_{x}-\nu_{x}}{\nu_{x}}-\log(\nu_{x})\right)\\ &=(\mu_{x}-\nu_{x})\cdot\frac{\mu_{x}}{\nu_{x}},\end{split}$ (3.21)
@@
This is a direct consequence of Lemma 3.9. Indeed, for $\varepsilon>0$ small enough we have
$\lVert\mu-\nu\rVert_{\infty}\leq\frac{2-\varepsilon}{2+\varepsilon}\cdot\min\{\mu_{x}:\mu_{x}>0\}$
and thus by Lemma 3.9 with $c=1+\varepsilon$ we have
$D_{\operatorname{KL}}(\mu,\nu)\leq(1+\varepsilon)\lVert\mu-\nu\rVert_{\operatorname*{\operatorname{TV}}}.$
As this holds for all $\varepsilon>0$ small enough, the claim follows. ∎
Before we prove the convergence of Fisher-Rao gradient flows we restate the result.
See 3.2
Proof.

### v2

@@ 1. Introduction
Natural policy gradient (NPG) methods and their proximal and trust region formulations known as PPO and TRPO are among the most popular policy optimization techniques in modern reinforcement learning (RL). As such they serve as a cornerstone of many recent RL success stories including celebrated advancements in computer games [51, 52, 11] and the recent development of large language models like ChatGPT [1]. This has motivated a quickly growing body of work studying the theoretical aspects such as the convergence properties and statistical efficacy of natural policy gradient methods. Almost all of these works consider a specific model geometry where the Fisher-Rao metrics of the individual rows of the policy are mixed according to their state distribution or slight modifications of this [26, 9, 36, 28]. However, other choices for the model geometry are possible. In particular, the Fisher metric on the state-action distributions has been used to design a natural gradient method as well as actor-critic and a trust-region variant known as relative entropy search (REPS) [37, 38, 45]. This alternative natural policy gradient has been found to have the potential to reduce the severity of plateaus [37] and improve the performance of actor-critic methods [38]. Despite these findings, theoretical results remain scarce. Initial works on the convergence of state-action natural policy gradients show an exponential convergence guarantee [42], without quantifying the exponential rate and without addressing function approximation. In this article, we provide quantitative convergence results for state-action natural policy gradients both with and without function approximation.
For our theoretical analysis, we work in the space of state-action distributions which brings the benefit that the reward optimization problem becomes a linear program [27]. In particular, for rich enough parametric policy models, the state-action natural policy gradient methods can be described by the Fisher-Rao gradient flow of the state-action linear program [42]. This motivates us to study Fisher-Rao gradient flows for general linear programs. These flows coincide with the solutions of entropy-regularized linear programs and thus by studying the convergence of the flow we also bound the error introduced by entropic regularization in linear programming.
@@ 1.1. Contributions
It is the goal of this article to provide insights into the convergence properties of state-action natural policy gradients with and without function approximation. To this end, we first provide an explicit convergence analysis of Fisher-Rao gradient flows of general linear programs. More precisely, our main contributions can be summarized as follows:
  • We study Fisher-Rao flows of general linear programs. Leveraging a local generalized strong convexity condition we show linear convergence both in KL-divergence and function value with an exponential rate depending on the geometry of the linear program, see Theorem 3.2 and Theorem 3.13 for unique and non-unique optimizers.
@@
  • We study natural gradients for parametric measures and show sublinear convergence under inexact gradient evaluations up to an approximation error and a distribution mismatch measured in the $\chi^{2}$-divergence, see Corollary 4.7.
@@
  • For non-unique optimizers, the asymptotic limit of Fisher-Rao gradient flows is known to be the information projection of the initial condition to the set of optimizers. We strengthen this by providing an exponential convergence rate, see Theorem 3.13, and by extending this result to state-action natural policy gradients. This shows that state-action natural gradients converge to an optimal policy that achieves maximal entropy over states and actions, which characterizes its implicit bias, see Theorem 3.13 and Corollary 5.8.
@@ 1.2. Related works
State-action natural policy gradients were recently studied with and without state-action entropy regularization in [42]. For regularization strength $\lambda>0$ that work showed $O(e^{-\lambda t})$ convergence, but in the unregularized case, the precise exponential rate was not characterized.
@@
In the case of a linear program, the Fisher-Rao gradient flow parametrized by time corresponds to the trajectory of solutions of the entropy-regularized program parametrized by the inverse regularization strength, which has been studied in several works. An exponential convergence result was obtained in [19] and subsequently, the rate was characterized as $O(e^{-\delta t})$ for a constant $\delta$ depending on the linear program [60, 54]. The results obtained in this article follow an alternative proof strategy and provide exponential convergence $O(e^{-\Delta t})$, where $\Delta\geq\delta$, where we show that the improvement can be arbitrarily large, see Example 3.5. This improvement can be strict for the linear programs encountered in Markov decision processes under standard assumptions. Whereas existing works study convergence in function value, our results also cover convergence in the KL-divergence. Finally, the geometry of Fisher-Rao gradient flows or equivalently the entropic central path was recently described as the intersection of the feasible region with a toric variety [53].
@@ 1.3. Notation and terminology
Finally, for a differentiable curve $(c_{t})_{t\in I}\subseteq\mathcal{M}$ defined on an interval $I\subseteq\mathbb{R}$ mapping to a manifold $\mathcal{M}$ we denote its time derivative by $\partial_{t}c_{t}$.
@@ 2. Preliminaries on Fisher-Rao Gradient Flows
To gain insight into natural gradient descent methods, we study their time-continuous version which is given by $\partial_{t}\theta_{t}=-F(\theta_{t})^{+}\nabla f(\mu_{\theta_{t}})$, where $\mu_{\theta}$ is a parametrized measure model and $f(\mu)$ is an objective function and $F(\theta)$ is the Fisher-information matrix [6]. The objective function can be a log-likelihood in the case of maximum likelihood estimation or a linear function in the case of reinforcement learning as we will see in Section 5. The Fisher-information matrix is closely connected to a specific Riemannian geometry, the Fisher-Rao metric, on the space of probability measures, which we introduce and discuss here. As we study gradient-based optimizers, we put a special emphasis on gradient flows with respect to the Fisher-Rao metric and provide a self-contained review of the properties of Fisher-Rao gradient flows that we require later. The results in this section can be generalized to a large class of Hessian geometries and – apart from the central path property – also to other objectives albeit with different proofs, for which we refer to [5, 39].
@@
where we denote the induced norm by $\lVert v\rVert_{g_{\mu}^{\operatorname*{\operatorname{FR}}}}\coloneqq{g_{\mu}^{\operatorname*{\operatorname{FR}}}(v,v)}^{\frac{1}{2}}$. The Fisher-Rao metric was introduced in the seminal works of C. R. Rao [48, 49] to provide lower bounds on the statistical error in parameter estimation known as the Cramer-Rao bound. This geometric approach to statistical estimation has subsequently led to the development of the field of information geometry, where N. N. Čencov characterized the Fisher-Rao metric as the unique Riemannian metric (up to scaling) that is invariant under sufficient statistics [18, 7, 8]. Despite its central role in statistics, our main motivation for studying the Fisher-Rao metric is for its use in reinforcement learning, where it has been used to design natural gradient algorithms as well as trust region methods [6, 37, 45]. Further, it is very closely related to entropic regularization in linear programming, which enjoys immense popularity, particularly in computational optimal transport [47, 54], see also [60] for a detailed discussion of entropy regularized linear programming.
@@
as it is induced by the Hessian of the (negative) entropy, meaning that we have $g_{\mu}^{\operatorname*{\operatorname{FR}}}(v,w)=v^{\top}\nabla^{2}\phi(\mu)w$ for all $v,w\in\mathbb{R}^{\mathbb{X}},\mu\in\mathbb{R}^{\mathbb{X}}_{>0}$. As such, the Fisher-Rao metric falls into the class of Hessian metrics that have been studied in convex optimization; we refer to [5, 39] for general well-posedness and convergence results. An important concept in the analysis of Hessian gradient flows is the Bregman divergence induced by $\phi$, which in the case of the negative entropy is given by the KL-divergence
@@
Proposition 2.2 (Central path property,[5]).
@@
On the other hand, for the gradient flow, we can use (2.7) and compute for $v\in TP$
$\langle\nabla\phi(\mu_{t})-\nabla\phi(\mu_{0}),v\rangle$ $=\int_{0}^{t}\partial_{s}\langle\nabla\phi(\mu_{s}),v\rangle{\mathop{}\!\mathrm{d}}s=\int_{0}^{t}\langle\nabla^{2}\phi(\mu_{s})\partial_{s}\mu_{s},v\rangle{\mathop{}\!\mathrm{d}}s$
$=\int_{0}^{t}\langle\nabla f(\mu_{s})\mu_{s},v\rangle{\mathop{}\!\mathrm{d}}s=\int_{0}^{t}\langle c,v\rangle{\mathop{}\!\mathrm{d}}s=t\langle c,v\rangle.$
@@
Corollary 2.3 (Sublinear convergence rate,[5]).
Consider Setting 2.1 and assume that the linear program (2.8) admits a solution $\mu^{\star}\in P$. Then for $\mu\in P$ it holds that
@@
We have $c^{\top}\mu_{t}-t^{-1}D_{\operatorname{KL}}(\mu_{t},\mu)\geq c^{\top}\mu^{\star}-t^{-1}D_{\operatorname{KL}}(\mu^{\star},\mu)$ by the central path property. Rearranging yields the result. ∎
@@
Theorem 2.4 (Well-posedness of FR GFs,[5]).
@@
The local existence and uniqueness follow from the Picard-Lindelöf theorem [56]. Hence, it suffices to show that the Fisher-Rao gradient flow does not hit the boundary $\partial P$ in finite time. By the central path property, this is equivalent to the statement that the solutions of all KL-regularized problems (2.9) lie in the interior $\operatorname{int}(P)$ of the polyhedron, which can be easily checked. ∎
@@ 3. Convergence of Fisher-Rao Gradient Flows
We have seen that Fisher-Rao gradient flows converge globally at a sublinear rate $O(t^{-1})$. We now build on this analysis and show that once the gradient flow enters a vicinity of the optimizer, it converges at a quasi-linear rate $O(t^{\kappa}e^{-\Delta t})$, where $\Delta>0$ depends on the geometry of the linear program and $\kappa>0$ depends on the initial condition $\mu_{0}$. Note that this yields $O(e^{-ct})$ convergence for all $c<\Delta$ and hence we also simply talk of a linear convergence rate. We consider linear programs of the following form.
@@
The following result is the main contribution of this article, where we defer the proof to Section 3.1. We first establish it under the assumption that the linear program (3.1) admits a unique solution and provide a generalization in Theorem 3.13.
@@
where $R_{H}\coloneqq\max_{\mu\in P}H(\mu)-\min_{\mu\in P}H(\mu)\leq\log\lvert\mathbb{X}\rvert$ denotes the entropic radius of $P$ and $\Delta>0$ and $t_{0}\geq 0$ are defined in (3.2) and (3.3), respectively.
@@
showing that Theorem 3.2 offers an improvement of these previous results.
For the special case $P=\Delta_{\mathbb{X}}$, for which a matching lower bound was constructed in [60], the two constants agree. More generally, it is easily checked that $\delta=\Delta$ if and only if there is a neighboring vertex $\mu\in N(\mu^{\star})$ which has minimal optimality gap $c^{\top}\mu^{\star}-c^{\top}\mu$ and has disjoint support from $\mu^{\star}$. To see this, note that for two probability vectors $\mu_{1},\mu_{2}\in\Delta_{\mathbb{X}}$ we have $\lVert\mu_{1}-\mu_{2}\rVert_{\textup{TV}}=\frac{1}{2}\lVert\mu_{1}-\mu_{2}\rVert_{1}\leq 1$ with $\lVert\mu_{1}-\mu_{2}\rVert=1$ if and only if $\mu_{1}$ and $\mu_{2}$ have disjoint support, meaning
$\{x\in\mathbb{X}:\mu_{1}(x)>0\}\cap\{x\in\mathbb{X}:\mu_{2}(x)>0\}=\emptyset.$
Hence, for $\mu\in N(\mu^{\star})$ without disjoint support from $\mu^{\star}$ we have $\lVert\mu^{\star}-\mu\rVert_{\textup{TV}}<1$. This implies that $\delta=\Delta$ if and only if there is a neighboring vertex $\mu\in N(\mu^{\star})$ which has minimal optimality gap $c^{\top}\mu^{\star}-c^{\top}\mu$ and has disjoint support from $\mu^{\star}$.
The constant $\Delta$ depends on the slope of $c$ along the outgoing edges and thus the local geometry of the feasible region around $\mu^{\star}$, where $\delta$ is simply based on the suboptimality at the neighboring vertices. Because of this, the difference between $\delta$ and $\Delta$ can be arbitrarily big as we show in Example 3.5. Further, for Markov decision processes the feasible region of the (dual) linear program is a strict subset $\mathscrsfs{D}\subsetneq\Delta_{\mathbb{S}\times\mathbb{A}}$ and under the standard exploratory 5.2 and more than one state we have $\delta<\Delta$, see Remark 5.11. In Section 5.2 we provide an explicit example of a Markov decision process where $\delta<\Delta$.
Further, for gradient flows with respect to a Riemannian metric of the form $g_{\mu}^{\sigma}(v,w)\coloneqq\sum_{x\in\mathbb{X}}\frac{v_{x}w_{x}}{\mu_{x}^{\sigma}}$ one can show $O(t^{-\frac{1}{\sigma-1}})$ convergence for $\sigma\in(1,2)$, see [42]. Note that this can be extended to the case $\sigma=2$, corresponding to logarithmic barriers for which the central path converges at a $O(t^{-1})$ rate [15, Section 11.2].
Example 3.5 (Arbitrarily large improvement).
We consider $\mathbb{X}=\{1,2,3,4\}$ and =$\mathcal{L}=\{\mu\in\mathbb{R}^{\mathbb{X}}:\mu(1)=\alpha\}$ for $\alpha\in(0,1)$. Then, the vertices of $P=\Delta_{\mathbb{X}}\cap\mathcal{L}$ are given by $(1-\alpha)\delta_{2},(1-\alpha)\delta_{3}$ and $(1-\alpha)\delta_{4}$, where $\delta_{i}$ denotes the Dirac at $i$. When choosing the cost $c=\delta_{2}$ we have $\delta=1-\alpha$ but $\Delta=1$. For $\alpha\nearrow 1$ the rate $\delta$ deteriorates towards 00, whereas $\Delta$ remains constant. The reason for this is that $\Delta$ depends on the slope of $c$ relative to the outgoing edges, whereas $\delta$ depends on the suboptimality of the neighboring vertices. Hence, $\delta$ can be smaller than $\Delta$ by an arbitrarily large factor.
@@ 3.1. Convergence of Fisher-Rao Gradient Flows
Lemma 3.8 ([5]).
Consider Setting 2.1, whereby we allow $f\colon\mathbb{R}_{>0}^{\mathbb{X}}\to\mathbb{R}$ to be an arbitrary differentiable function, and fix $\mu\in P$. Then for any $t\geq 0$, it holds that
@@
Using (3.13) and (3.10) we find that for all $t\geq T$ it holds that $\partial_{t}D_{\operatorname{KL}}(\mu^{\star},\mu_{t})=c^{\top}\mu_{t}-c^{\top}\mu^{\star}\leq-\kappa_{t}D_{\operatorname{KL}}(\mu^{\star},\mu_{t})$. Now Gronwall’s inequality yields (3.11). By Corollary 2.3 we have for any $h>0$ that
$c^{\top}\mu^{\star}-c^{\top}\mu_{t}\leq\frac{D_{\operatorname{KL}}(\mu^{\star},\mu_{t-h})}{h}\leq D_{\operatorname{KL}}(\mu^{\star},\mu_{0})\cdot\frac{\exp\left(-\int_{t_{0}}^{t-h}\kappa_{s}{\mathop{}\!\mathrm{d}}s\right)}{h}.$
@@
and further
@@
If $F^{\star}\neq P$, then $c^{\top}\mu^{\star}-c^{\top}\mu>0$ for some vertex $\mu$, which implies $\Delta>0$. To simplify notation we denote the set $E\coloneqq\{\mu-\mu^{\star}:\mu\in N(\mu^{\star})\setminus F^{\star},\mu^{\star}\in\operatorname{vert}(F^{\star})\}$ of edges such that exactly one of the two endpoints is contained in $F^{\star}$. Then, the polytope $P$ is contained in
see Lemma A.1 and hence we can write $\mu\in P$ as $\mu=\mu^{\star}+\sum_{e}\alpha_{e}e$ for some $\mu^{\star}\in F^{\star}$. Using the triangle inequality we obtain
$\Delta\lVert\mu^{\star}-\mu\rVert\leq\Delta\sum_{e\in E}\alpha_{e}\lVert e\rVert\leq-\sum_{e\in E}\alpha_{e}c^{\top}e=c^{\top}\mu^{\star}-c^{\top}\mu.$
@@
Consider a finite set $\mathbb{X}$ and a probability distribution $\mu\in\Delta_{\mathbb{X}}$. Let $c>1$ and set $\delta\coloneqq\frac{c-1}{c+1}\cdot\min\{\mu_{x}:\mu_{x}>0\}>0$. Then for all $\nu\in\Delta_{\mathbb{X}}$ satisfying $\lVert\mu-\nu\rVert_{\infty}\leq\delta$ it holds that
@@
$\begin{split}\mu_{x}\log\left(\frac{\mu_{x}}{\nu_{x}}\right)&=\mu_{x}\left(\log(\nu_{x}+(\mu_{x}-\nu_{x}))-\log(\nu_{x})\right)\\ &\leq\mu_{x}\left(\log(\nu_{x})+\frac{\mu_{x}-\nu_{x}}{\nu_{x}}-\log(\nu_{x})\right)=(\mu_{x}-\nu_{x})\cdot\frac{\mu_{x}}{\nu_{x}},\end{split}$ (3.17)
@@
This is a direct consequence of Lemma 3.10. Indeed, for $\varepsilon>0$ small enough we have $\lVert\mu-\nu\rVert_{\infty}\leq\frac{2-\varepsilon}{2+\varepsilon}\cdot\min\{\mu_{x}:\mu_{x}>0\}$ and thus by Lemma 3.10 with $c=1+\varepsilon$ we have $D_{\operatorname{KL}}(\mu,\nu)\leq(1+\varepsilon)\lVert\mu-\nu\rVert_{\operatorname*{\operatorname{TV}}}$. Note that $\varepsilon>0$ was arbitrary. ∎
Now we prove our main result on the convergence of Fisher-Rao gradient flows.
Proof of Theorem 3.2.